OpenStudy (anonymous):

What are the solutions? 1/2x^2+2x+3=0

OpenStudy (shamil98):

Use the quadratic formula.

OpenStudy (shamil98):

\[\huge ax^2 + bx + c = 0\] \[\huge x = \frac{ -b \pm \sqrt {b^2 -4ac} }{ ? }\]

OpenStudy (shamil98):

the bottom is 2a*

OpenStudy (loser66):

1/ is order of the problem or 1 is 1/.........

OpenStudy (ddcamp):

Hmm, I don't think that that has any real solutions. Have you started working with complex numbers yet?

OpenStudy (anonymous):

Yes @DDCamp but I'm really bad at understanding things, so it takes me a while to actually get it

OpenStudy (ddcamp):

And, on the first term, is it (1/2)x² or 1/(2x²)?

OpenStudy (anonymous):

(1/2)x²

OpenStudy (ddcamp):

OK. In that case, just use the quadratic equation like @shamil98 said. Show me how far you can simplify it when you've plugged the numbers in.

OpenStudy (anonymous):

I'm pretty sure I have it right so far.